the official volleyball measures 27 inches in circumference. what is the surface area?
Answer by checkley71(8403)

TSA=4PIR^2
TSA=4*3.14(27/2)^2
TSA=12.56(13.5)^2
TSA=12.56*182.25
TSA=2,289.06 SQUARE INCHES IS THE TOTAL SURFACE AREA.